Solve the following equation involving fractions and variables
-4/9(2X-4)=48
Please explian how to multiply a fraction when variables are present
-4/9 (2X) -4/9 (-4) = 48
Help me please!

Answer by ankor@dixie-net.com(6693)

Assume the problem is:

Get rid of the denominator, multiply both sides by 9, results:
-4(2x - 4) = 9(48)
Multiply what's inside the brackets by -4, results:
-8x + 16 = 432
:
-8x = 432 - 16
:
-8x = 416
x has to be positive, multiply both sides by -1
8x = -416
Divide both sides by 8
x =
x = -52
;
:
Check solution in original equation, replace x with -52



Cancel 9 into 108
-4(-12) = 48
+48 = 48; confirms our solution
:
:
:
Did you understand this?